OpenStudy (anonymous):

The radius of a circular disk is given as 28 cm with a maximum error in measurement of .4 cm. Use differentials to estimate the maximum possible error in the calculated area of the disk. What is the relative error?

OpenStudy (campbell_st):

well the maximum radius would be 28.4 cm minumum radius would be 27.6 cm use theese values to find the difference in the areas.

OpenStudy (anonymous):

which would be 140.7433509 so where do i go from there

OpenStudy (campbell_st):

well isn't relative error the absolute error divided by the measurement expressed as a percentage. so what is the area when r = 28 and perhaps relative error is 140. 7/(Area) * 100 thats my best guess

OpenStudy (agent0smith):

From \[A=πr^2 \] find dA/dr: \[\frac{dA}{dr}=2πr\] so \[dA=2πrdr \] where dr is the error in the radius and so dA is the error in the area. dr will be 0.4 since the error is plus or minus 0.4cm.

OpenStudy (agent0smith):

And r is the radius, 28cm. To find relative error, just find: \[\frac{ dA }{A} = ?\] and \[A = \pi r^2 \] (where r=28cm)